OpenStudy (anonymous):

help help !!!!! y varies jointly as x and z. find y for the given values of x and z. y =32 when x=8 and z=2; x=4 and z=-2

OpenStudy (aroub):

y=kxz 32=k(8)(2) 32=k(16) k=2 y=kxz y=2(4)(-2) y=-16